# Adversarial Poetry as a Universal Single-Turn Jailbreak Mechanism in Large Language Models

### v1

@@ 6.6 Future Works
This study highlights a systematic vulnerability class arising from stylistic distribution shifts, but several areas require further investigation. First, we plan to expand mechanistic analysis of poetic prompts, including probing internal representations, tracing activation pathways, and isolating whether failures originate in semantic routing, safety-layer heuristics, or decoding-time filters. Second, we will broaden the linguistic scope beyond English to evaluate whether poetic structure interacts differently with language-specific training regimes. Third, we intend to explore a wider family of stylistic operators – narrative, archaic, bureaucratic, or surrealist forms – to determine whether poetry is a particularly adversarial subspace or part of a broader stylistic vulnerability manifold. Finally, we aim to analyse architectural and provider-level disparities to understand why some systems degrade less than others, and whether robustness correlates with model size, safety-stack design, or training data curation. These extensions will help clarify the boundaries of stylistic jailbreaks and inform the development of evaluation methods that better capture generalisation under real-world input variability.

### v2

@@ 6.6 Future Works
This study highlights a systematic vulnerability class arising from stylistic distribution shifts, but several areas require further investigation.
First, we plan to expand mechanistic analysis of poetic prompts, including probing internal representations, tracing activation pathways, and isolating whether failures originate in semantic routing, safety-layer heuristics, or decoding-time filters.
Second, we will broaden the linguistic scope beyond English to evaluate whether poetic structure interacts differently with language-specific training regimes. Third, we intend to explore a wider family of stylistic operators – narrative, archaic, bureaucratic, or surrealist forms – to determine whether poetry is a particularly adversarial subspace or part of a broader stylistic vulnerability manifold.
Finally, we aim to analyse architectural and provider-level disparities to understand why some systems degrade less than others, and whether robustness correlates with model size, safety-stack design, or training data curation. These extensions will help clarify the boundaries of stylistic jailbreaks and inform the development of evaluation methods that better capture generalisation under real-world input variability.